OpenStudy (anonymous):

Help pleaseee! The height of water shooting from a fountain is modeled by the function f(x) = −4x2 + 24x − 29 where x is the distance from the spout in feet. Complete the square to determine the maximum height of the path of the water. −4(x − 3)2 − 29; The maximum height of the water is 3 feet. −4(x − 3)2 − 29; The maximum height of the water is 29 feet. −4(x − 3)2 + 7; The maximum height of the water is 7 feet. −4(x − 3)2 + 7; The maximum height of the water is 3 feet.

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

chitty chitty @bangbang559

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

http://www.imdb.com/title/tt0062803/

OpenStudy (anonymous):

its C

OpenStudy (anonymous):

how do you know?

OpenStudy (anonymous):

@hoblos

OpenStudy (anonymous):

@zepdrix @ganeshie8

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Come on back, and I'll work the problem with you...

OpenStudy (anonymous):

here

OpenStudy (whpalmer4):

\[f(x) =−4x^2 + 24x − 29\]We need to complete the square. I would start by factoring out the -4: \[-4(x^2-6x-\frac{29}{4})\]Do you agree with that result so far?

OpenStudy (whpalmer4):

Actually, let me write it slightly differently: \[f(x) = -4(x^2-6x) -29\]

OpenStudy (anonymous):

yup

OpenStudy (whpalmer4):

Okay. Now to complete the square, we take half of the coefficient of \(x\), and square it. \[(\frac{-6}{2})^2 = (-3)^2 = 9\]We are going to simultaneously add and subtract that inside the parentheses. \(9-9=0\) so we are not changing the value of anything. \[f(x) = -4(x^2 -6x +9 - 9) - 29\]Now, we can rewrite \[(x^2-6x+9) = (x-3)(x-3)\]\[f(x) = -4((x-3)^2 -9) - 29\]We apply the distributive property: \[f(x) = -4(x-3)^2 -4(-9) - 29\]\[f(x) = -4(x-3)^2+36-29\]\[f(x = -4(x-3)^2+7\]Any questions about what I did?

OpenStudy (whpalmer4):

That last line should be \[f(x) =-4(x-3)^2+7\]

OpenStudy (whpalmer4):

Vertex form for a parabola is \[y = a(x-h)^2+k\]where the vertex is located at \((h,k)\) Can you tell me the vertex of this parabola?

OpenStudy (anonymous):

this really confuses me

OpenStudy (whpalmer4):

What does, the completing of the square, or vertex form, or both?

OpenStudy (anonymous):

a new equation

OpenStudy (whpalmer4):

which new equation are you talking about?

OpenStudy (whpalmer4):

\[f(x) = -4(x-3)^2 + 7\]That's the equation we got by completing the square. \[y = a(x-h)^2 + k\]That's the formula of a parabola in "vertex" form. Notice the similarity? The two are equal if we say \[y = f(x)\]\[a=-4\]\[h=3\]\[k=7\]Therefore, our parabola has a vertex of \((3,7)\). As \(a < 0\), that parabola opens downward, and the vertex will be at the highest point. Therefore, the maximum height of the water is 7 feet, which is the y value of the vertex.

OpenStudy (whpalmer4):

A plot of the function is attached, where you can clearly see that the vertex is at (3,7).